APPRENTICE and SIT-INS

Name (34)  @uchicago.edu
Robert Bishop  rebishop
Siddhartha Biswas  sbiswas
Isaac Davis  idavis
Kristin Dean  kristindean
Marcello Delgado  madelgado
Wai Lee Chin Feman  wchinfeman
Madelyn Freed  madelyn
Namita Gupta  namita1025
Hong Gyun Kim  diminate
Korei Klein  KoreiKlein
Kegon Tan Teng Kok  kegontantk
Maxwell Levine  mslevine
Chenchuan Li  chenchuan
Nikhil Maddirala  nikhilm
James Murphy  murphyj1
Jia Hong Ray Ng  raynjh
Ben Nordstrom  bnordstrom
Stephen Patrias  cupioarcanamundi
Aaron Plavnick  aplavnick
James Salvatore  jsal13
Mira Scarvalone  mira.scarvalone@gmail.com
Nathan Schulz  nathanschulz
Jeremy Schwartz  Jschwartz
Joohyun Shon  shon
Alexander Staples-Moore  asm
Matthew Steffen  matthew.steffen@gmail.com
Caroline Terry  caterry
Justin Walker  justinw
Issac Ottoni Wilhelm  spaceka13
John Wiltshire-Gordon  jhdwg
Angelica Wong  angelicawong
Xiaochen Wu  xwu6
Karen J. Ye  kye
Richard Yeung  richardyeung

SESAME

Name (23)  @uchicago.edu
William C. Abram  abramwc
Dylan Allegretti  allegretti
Ari Allyn-Feuer  aallynfeuer
Elan Bechor  ebechor
Sarah Bennett  sbennett
John Binder  JBinder
Beau Dabbs  bdabbs13
Joseph Doliner  jdoliner
Jonathan Emberton  jemberton
Daniel Feltey  dfeltey
Patrick Greene  patrickg
Aaron Halper  ahalper
Justin Huang  jhuang4
Ethan Jerzak  jerzak
Dennis Kriventsov  krivents
Carmel Levy  carmel1
Brian Libgober  libgober
Thomas Milnor  Heahengel@gmail.com
Alexander J. Munk  amunk
Peter Nelson  petern
Matt Olson  molson2
Alex Rosenfeld  arosenfe
Michael Travis  mtravis

COLLEGIATE SCHOLARS

Name (4)  @uchicago.edu
Hikaru Michael Kiyo  mkiyo3
Guillaume Pouliot  guillaumepouliot
Seung Hyun Yang  yang.seunghyun@gmail.com
Tengren Zhang  tengren85

YSP: 7–8

Name (8)  @uchicago.edu
Ioana Oriana Bercea  ioana
Michael D. Calderbank MCalderbank
Lauren Cote  Lecote
Dustin Hedmark  dhedmark
Chris Henderson  ckhenderson
Alice Mark  amark
Ellen McCullagh  ellenm
Brooke Ullery  bullery

YSP: 9–10

Name (7)  @uchicago.edu
Jahnavi Bhaskar  jbhaskar
Patricia Brent  pmb
Eleanor Brush  ebrush
John Dyer  dyer
Alex Fix  afix
Ariel Hafftka  ahafftka
Noah Schweber  schweber

YSP: 11–12

Name (5)  @uchicago.edu
Corrin J. Clarkson  clarkson
Nicholas Ford  nford
Prerna Nadathur  pnadathur
Caitlin Waddle  kwaddle
Matthew Woolf  mwoolf